OpenStudy (anonymous):

The rectangular prism shown has a height of , a length of , and a width of . Part A: Suppose you have cubes that are on each side. How many of these cubes would it take to fill a cube that is 1 inch high, 1 inch long, and 1 inch wide? Part B: Use your answer to Part A to find the volume of one of the cubes. Part C: How many cubes would it take to fill the prism shown? Part D: Use your answers to Parts B and C to find the volume of the prism. Part E: Multiply the height, length, and width of the prism. Compare this product to your answer to Part D.

OpenStudy (anonymous):

height of 2 1/2 inches, a length 1 1/2 inches and a width of 1/2 inch

OpenStudy (anonymous):

V of rectangular prism is l*w*h and since your cubes are just 1 inch^3 your answer for a will just be V

OpenStudy (anonymous):

For B a small cube would just be 1 inch cuped

OpenStudy (anonymous):

IC TY:)

OpenStudy (anonymous):

I have another, these are just confusing to me

OpenStudy (anonymous):

You need to provide more pictures and graphs. I am just assuming based on what you typed

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Mya has a raised garden bed in the shape of a rectangular prism. The bed is 8 feet long, 3 2/3 feet wide, and 2 feet high. Part A: Belinda wants to make a raised garden bed that covers the same area as Mya’s, but the space she has is only 3 feet wide. How long should her garden bed be? Part B: Find the volume of Mya’s gardening bed. Suppose Belinda’s garden bed is the same height as Mya’s. Will Belinda’s bed have the same volume? Explain. Part C: A small bag of soil contains 2 cubic feet. How many bags of soil should Mya buy if she wants to fill her garden bed completely? Show your work. Part D: Belinda already has 22 bags of soil. Use your answer to Part C to find the depth to which she can fill her garden bed.

OpenStudy (anonymous):

well mya's is 8*(3+2/3)=A since belinda wants same area but is limited to 3 feet wide then X*(3)=8*(3+2/3) that is part A

OpenStudy (anonymous):

solve for x

OpenStudy (anonymous):

For part B the volume is just l*w*h another way to express volume is area * h since both bel and myas gardens are same area and same height then they should have same volume.

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

For C just divide what you got in Part b by 2 garden volume/ 2 which is the soils volume

OpenStudy (anonymous):

for part D 22 bags of soil * 2 is the volume of the soil so you have a total volume of 44 so 44= w*l*h 44= 8*(3+2/3) * h solve for h

OpenStudy (anonymous):

i got 29.28 for bel, is that right

OpenStudy (anonymous):

28.28 as volume?

OpenStudy (anonymous):

area

OpenStudy (anonymous):

they should both have same area

OpenStudy (anonymous):

I guess I'm having trouble with the length

OpenStudy (anonymous):

29.33333334 is area

OpenStudy (anonymous):

so 29.33333334=x*3 just divide by 3 and you get length 9.777777780

OpenStudy (anonymous):

I was way off, oh i get it now

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

i have 42.66 for v

OpenStudy (anonymous):

no :( 58.66666667 8*(3+2/3)*2.

OpenStudy (anonymous):

I'm glad you're on tonight:)

OpenStudy (anonymous):

remember that V=l*w*h and you can multiply in any order you want. in othher words whl, hlw, whl... that helps a lot

OpenStudy (anonymous):

I have another assignment finished, would you be willing to make sure it's ok? I can message it to you

OpenStudy (anonymous):

no real rush though

OpenStudy (anonymous):

I can check your work and give you hints sure.

OpenStudy (anonymous):

k I'll send it as soon as I finish this other assignment